# Joint Resource-Power Allocation and UE Rank Selection in Multi-User MIMO Systems with Linear Transceivers

### v1

@@ IV-A Uplink Scheduling Information
Having obtained the UE transmit powers, the UE ranks, and the resource allocation indices, the BS must now transmit this information to each co-scheduled UE on a downlink control information (DCI) [15] message. However, the standard DCI formats in 5G NR can only encode a single low-resolution closed-loop power control command per UE and slot. Instead, the MU-MIMO power-control solution we propose requires transmit power control (TPC) commands that are PRBG-, slot-, and layer-specific. 5G NR DCI formats are therefore insufficient to convey these data-rich BS decisions. For this reason, we propose the following additional features to support MU-MIMO TPC commands. Note that these features differ in the amount of signaling overhead and consequently, not all of them may be preferred in all scenarios.
@@
    High granularity DCI: DCI encodes the discretized full set of transmit powers for each PRBG on each layer. This incurs the highest overhead and might be practical in 6G deployments with extreme bandwidths.
@@
    Medium granularity DCI: A condensed set of discretized transmit powers, only one per layer. In this case, the UE uses the same transmit power for all scheduled PRBGs per layer. This embodiment may yield the best trade-off between performance and signaling overhead, and it may be optimal in deployments with low-frequency selectivity, such as rural outdoors.
@@
    Low granularity DCI: A single discretized transmit power per UE. In this case, the UE uses a common transmit power for all the scheduled PRBs and layers. This incurs the least overhead, and depending on the case, might not significantly degrade the performance compared to the having separate transmit powers per PRBG and layer.
@@ IV-B Possible Limitations of the proposed technique
If some of the UEs have poor channel conditions (e.g., cell-edge UEs) and an insufficient power budget to ensure a certain SINR at the BS, the BS would not have a minimum required level of SINR to obtain good quality channel estimates from the SRS of these UEs. This scenario would adversely impact the downlink more because the assumed ZF precoder design is no longer valid.
@@ IV-C Computational Complexity and Practical Considerations
where $\mathbf{x}_{L}\in[-\infty,\infty)^{n}$, $\mathbf{x}_{U}\in(-\infty,\infty]^{n}$ with $\left[\mathbf{x}_{L}\right]_{i}<\left[\mathbf{x}_{U}\right]_{i}$, $\forall i=1,\cdots,n$, $\mathbf{c}_{L}\in[-\infty,\infty)^{m}$, $\mathbf{c}_{U}\in(-\infty,\infty]^{m}$ with $\left[\mathbf{c}_{L}\right]_{i}<\left[\mathbf{c}_{U}\right]_{i}$, $\forall i=1,\cdots,m$. The objective function is $f:\mathbb{R}^{n}\to\mathbb{R}$ and the constraint function is $c:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with any equality constraint being set by choosing $\mathbf{c}_{L}=\mathbf{c}_{U}$). For a tolerance $\epsilon>0$, the number of required iterations to obtain a solution is ${\cal O}\left(\sqrt{n}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$ [16]. Further, each iterations involves solving a linear system of equations that has a worst-case complexity of ${\cal O}\left(n^{3}\right)$. Note that this is an upper bound and dependent on the form of the objective function and how entangled the variables are with one another, and most practical linear solvers have a complexity less than ${\cal O}\left(n^{3}\right)$. Nevertheless, the worst-case complexity of any interior-point method in order to achieve a tolerance of $\epsilon>0$ is ${\cal O}\left(n^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$.
@@
The proposed technique uses two optimization stages with each optimization stage solving a convex optimization problem with $n=n_{RBG}n_{U}$ variables where $n_{U}=\sum_{i}n_{U}^{(i)}$. For the uplink problem (18a), the nature of the objective and the constraints imply that the optimization for each UE can be done separately. So, this translates to solving $N_{UE}$ independent optimization problems with $n=n_{RBG}n_{U}^{(i)}$ variables for UE $i$ in the first stage and $n=n_{RBG}n_{l}^{(i)}$ variables in the second stage. But each stage of the proposed technique requires the computation of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$, $\forall g=1,\cdots,n_{RBG}$, where $\hat{\mathbf{H}}_{g}\in\mathbb{C}^{n_{B}\times n_{U}}$, which has a worst-case complexity of ${\cal O}\left(n_{RBG}n_{U}^{3}\right)$. So, the worst-case complexity of the proposed technique for the uplink is ${\cal O}\left(\left(n_{RBG}n_{U,max}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)+n_{RBG}n_{U}^{3}\right)$, where $n_{U,max}=\max_{i\in{\mathcal{N}}_{UE}}\{n_{U}^{(i)}\}$ for an error tolerance of $\epsilon$.
@@
For the downlink, all the $n=n_{RBG}n_{U}$ variables need to be jointly optimized, leading to a worst-case complexity of ${\cal O}\left(\left(n_{RBG}n_{U}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$. Note that while the ZF precoders need to be calculated in order to obtain the $\lambda_{i,j,g}$, this is anyway unavoidable for downlink transmission and hence not part of the additional complexity resulting from the usage of the proposed technique. Further, note that since the Hessian of the objective function as shown in (19a) allows a block-diagonal structure, each iteration of the optimization problem involves solving a sparse linear system of equations which significantly reduces the overall complexity. If PRBG allocation is not considered, i.e., all the available PRBG are allocated to every UE, the complexity is significantly reduced from ${\cal O}\left(\left(n_{RBG}n_{U}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$ to ${\cal O}\left(n_{U}^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$, but this might come at the cost of throughput performance, especially in highly frequency selective channels.
@@
The computational burden of our method, requiring the execution of IPOPT every slot, presents a substantial challenge for real-time 6G implementation. To address this, we propose investigating both supervised and reinforcement learning approaches. Supervised learning could involve training a model on a dataset of optimal solutions, allowing for rapid prediction in subsequent slots. Alternatively, reinforcement learning, guided by our proposed technique as an expert policy, could uncover even better performing methods. Future research will focus on exploring these ML-based techniques, aiming to balance computational efficiency and performance.
@@ V Simulation Results
Maximum number of co-scheduled UEs per slot $8$
$n_{U}^{(i)}$ $4$ ($2$H×$\times 1$V×$\times 2$P), $\forall i=1,\cdots,8$
@@
Number of cells in the network $21$
Total number of UE in the network $210$
@@
Channel model 38.901 Urban Macro (UMa) NLoS [17, Section 7.2]
@@
UE OLPC parameters $(P_{0},\alpha)$ $P_{0}$ (dBm) $\in\{-85,-90,-100,-110\}$, $\alpha\in\{0.85,1\}$
Channel coding 5G NR LDPC [15, Section 5.3.2] with rate-matching
@@
MCS Table 5.1.3.1-2 (uplink) and Table 5.1.3.1-4 (downlink) [8]
@@
Hybrid automatic repeat request (HARQ) No
Number of independent UE drops 10
@@
Downlink precoder Block diagonalization ZF precoder [1]
@@
Number of iterations of IPOPT $10$
Figure 2: Comparison between the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot for the isolated cell case. The error bars denote the $90\%$ confidence interval.
Figure 3: Percentile ranks of the UE transmit power (left) and the PRB allocation for the isolated cell case.
  1. 1.
  2. 2.
    Cells with ICI from the adjacent cells whose statistics varies from slot to slot depending on the paired UEs in the adjacent cells. In this case, there is no way of obtaining an accurate enough estimate of $\mathbf{R}_{g}$ (introduced in Section III-A) at the BS in order to accurately estimate the correct MCS for all the co-scheduled UEs unless the ICI is lower than or comparable to the thermal noise level.
For the uplink, we use QPSK, 16-QAM, 64-QAM, and 256-QAM while for the downlink, we also use 1024-QAM. Explicit data transmission, receiver equalization, and channel coding/decoding using 5G NR LDPC codes are performed. We assume the availability of perfect CSI but no open loop link-adaptation (OLLA) [18] or HARQ which, in practice, can be expected to make up for ill-effects of imperfect CSI. Therefore, MCS selection is done as detailed in [19] while taking $\mathbf{R}_{g}$ to be $\sigma^{2}\mathbf{I}$ in (3), where $\sigma^{2}$ is the thermal noise variance per PRB.
V-A Uplink Results
Baseline power control: The total transmit power $P_{i}$ (dBm) for UE $i$ is obtained by setting the OLPC parameters $(P_{0},\alpha)$ in the following [20, Section 7]:
$P_{i}=\min\{P_{U,max},P_{0}+10\mathop{\mathrm{log}_{10}}\left(n_{PRB}\right)+\alpha PL\}$ (23)
[…]
Baseline PRBG allocation: We assume that all the PRBGs are allocated to the UE unless power-limited. In the latter case, we assign a number of PRBGs over which the UE has sufficient power to transmit, choosing the available PRBGs with the strongest channel gains for that UE. Then, the number of PRBGs allocated to the UE is taken to be
$n_{RBG,i}=\min\left\{n_{RBG},\max\left\{n_{RBG,min},\left\lfloor 10^{\frac{P_{U,max}-P_{0}-\alpha PL}{10}}\right\rfloor\right\}\right\}.$ (24)
@@ V-A Uplink Results
Figure 4: Comparison between the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot with ICI. The error bars denote the $90\%$ confidence interval.
Figure 5: Percentile ranks of the UE transmit power (left) and the PRB allocation (right) for the case with ICI.
Figure 6: Percentile ranks of the number of data layers per UE and the number of spatial layers per BS for the case with ICI.
Fig. 2 shows the comparison (“BL" indicates “baseline") of the average GM and AM UE rates, both computed per slot and then averaged across slots and independent UE drops. While the proposed technique is not targeted at maximizing the AM rates, we have shown them for illustration. Since there is no ICI, it is obvious that aggressively using as much of the available transmit power as required is generally the best strategy. Therefore, we use $r_{min}=0.23$ (corresponding to 4-QAM with the least coding rate in 5G NR) and $r_{max}=8$ (corresponding to the upper bound on the rate per data symbol for 256-QAM) for the proposed technique (Section III-A). While the proposed technique not only achieves the best GM UE rates, it also does so while requiring lower UE transmit powers than the naive, full transmit power scheme, as shown in Fig. 3.
For the case with ICI, since the transmit powers of the served UEs affect the reception in the neighbouring cells and the statistics of this ICI varies slot-to-slot, the best performing techniques usually choose the transmit powers conservatively so as to minimize the ICI. Otherwise, MCS selection would be significantly affected even for the baseline schemes. So, we conservatively take $r_{max}$ to be between $1$ and $2$. We also consider $\gamma=0.1$ for one of the baseline schemes so that a higher number of layers is selected for each UE. Fig. 4 highlights that the proposed scheme provides a significantly higher ($>50\%$) GM rates compared to the best baseline schemes while requiring a similar level of transmit powers per UE as shown in Fig. 5. The empirical cumulative distribution functions (CDFs) of the number of spatial layers per UE and the total number of spatial layers per BS are shown in Fig. 6. The trend clearly shows that while the AM rates can be improved by increasing $r_{max}$ at the cost of GM rates, the best balance is achieved for $r_{max}$ between $1.25$ and $1.75$.
Figure 7: Comparison between the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot. The error bars denote the $90\%$ confidence interval.
Figure 8: Percentile ranks of the number of data layers per UE and the number of spatial layers per BS.
Figure 9: Percentile ranks of the BS transmit power (left) and the UE PRB allocation (right).
@@ V-B Downlink Results
Baseline power allocation: Let $\mathbf{w}_{g,row}^{(k)}$ denote the $k^{th}$ row of $\mathbf{W}_{g}$, $k=1,\cdots,n_{B}$, $g=1,\cdots,n_{RBG}$, where $\mathbf{W}_{g}$ is the equivalent of $\mathbf{W}_{f,t}$ mentioned in Section II-B at a PRBG level. The columns of $\mathbf{W}_{g}$ are normalized to unity. Then, the precoder is taken to be $\mathbf{W}_{g}\leftarrow\sqrt{P}\mathbf{W}_{g}$ with $P=P_{ant}n_{RBG}/max_{k=1,\cdots,n_{B}}\left\{\sum_{g\in{\cal G}}n_{F}\|w_{g,row}^{(k)}\|^{2}\right\}$.
Baseline PRBG allocation: For the lack of a better alternative baseline scheme, we assume that all the PRBGs are allocated to all the UEs. Note that the techniques highlighted in [6][21] require reinforcement-learning and multiple PRB-looping which are out of scope of this paper.

### v2

@@ IV-A Uplink Scheduling Information
Having obtained the UE transmit powers, the UE ranks, and the resource allocation indices, the BS must now transmit this information to each co-scheduled UE on a downlink control information (DCI) [15] message. However, the standard DCI formats in 5G NR can only encode a single low-resolution closed-loop power control command per UE and slot. Instead, the MU-MIMO power-control solution we propose requires transmit power control (TPC) commands that are PRBG-, slot-, and layer-specific. 5G NR DCI formats are therefore insufficient to convey these data-rich BS decisions. For this reason, we propose the following additional features to support MU-MIMO TPC commands. Note that these features differ in the amount of signaling overhead, and consequently, not all of them may be preferred in all scenarios.
@@
    High granularity DCI: The DCI encodes the discretized full set of transmit powers for each PRBG on each layer. This incurs the highest overhead and might be practical in 6G deployments with extreme bandwidths.
@@
    Medium granularity DCI: A condensed set of discretized transmit powers, only one per layer, is sent. In this case, the UE uses the same transmit power for all scheduled PRBGs per layer. This embodiment may yield the best trade-off between performance and signaling overhead, and it may be optimal in deployments with low-frequency selectivity, such as rural outdoors.
@@
    Low granularity DCI: A single discretized transmit power per UE is transmitted on the DCI. In this case, the UE uses a common transmit power for all the scheduled PRBs and layers. This incurs the least overhead, and depending on the case, might not significantly degrade the performance compared to the having separate transmit powers per PRBG and layer.
@@ IV-B Possible Limitations of the proposed technique
If some of the UEs have poor channel conditions (e.g., cell-edge UEs) and an insufficient power budget to ensure a certain SINR at the BS, the BS would not be able to obtain good quality channel estimates from the SRSs of these UEs. This scenario would adversely impact the downlink more because the assumed ZF precoder design would no longer be valid.
@@ IV-C Computational Complexity and Practical Considerations
where $\mathbf{x}_{L}\in[-\infty,\infty)^{n}$, $\mathbf{x}_{U}\in(-\infty,\infty]^{n}$ with $\left[\mathbf{x}_{L}\right]_{i}<\left[\mathbf{x}_{U}\right]_{i}$, $\forall i=1,\cdots,n$, $\mathbf{c}_{L}\in[-\infty,\infty)^{m}$, $\mathbf{c}_{U}\in(-\infty,\infty]^{m}$ with $\left[\mathbf{c}_{L}\right]_{i}<\left[\mathbf{c}_{U}\right]_{i}$, $\forall i=1,\cdots,m$. The objective function is $f:\mathbb{R}^{n}\to\mathbb{R}$ and the constraint function is $c:\mathbb{R}^{n}\to\mathbb{R}^{m}$ (with any equality constraint being set by choosing $\mathbf{c}_{L}=\mathbf{c}_{U}$). For a tolerance $\epsilon>0$, the number of required iterations to obtain a solution is ${\cal O}\left(\sqrt{n}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$ [16]. Further, each iterations involves solving a linear system of equations that has a worst-case complexity of ${\cal O}\left(n^{3}\right)$. Note that this is an upper bound and dependent on the form of the objective function and how entangled the variables are with one another, and most practical linear solvers have a complexity less than ${\cal O}\left(n^{3}\right)$. Nevertheless, the worst-case complexity of any interior-point method in order to achieve a tolerance of $\epsilon>0$ is ${\cal O}\left(n^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$.
@@
The proposed technique uses two optimization stages with each optimization stage solving a convex optimization problem with $n=n_{RBG}n_{U}$ variables where $n_{U}=\sum_{i}n_{U}^{(i)}$. For the uplink problem (17a), the nature of the objective and the constraints imply that the optimization for each UE can be done separately. So, this translates to solving $N_{UE}$ independent optimization problems with $n=n_{RBG}n_{U}^{(i)}$ variables for UE $i$ in the first stage and $n=n_{RBG}n_{l}^{(i)}$ variables in the second stage. But each stage of the proposed technique requires the computation of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$, $\forall g=1,\cdots,n_{RBG}$, where $\hat{\mathbf{H}}_{g}\in\mathbb{C}^{n_{B}\times n_{U}}$, which has a worst-case complexity of ${\cal O}\left(n_{RBG}n_{U}^{3}\right)$. So, the worst-case complexity of the proposed technique for the uplink is ${\cal O}\left(N_{UE}\left(n_{RBG}n_{U,max}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)+n_{RBG}n_{U}^{3}\right)$, where $n_{U,max}=\max_{i\in{\mathcal{N}}_{UE}}\{n_{U}^{(i)}\}$.
@@
For the downlink, all the $n=n_{RBG}n_{U}$ variables need to be jointly optimized, leading to a worst-case complexity of ${\cal O}\left(\left(n_{RBG}n_{U}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$. Note that while the ZF precoders need to be calculated in order to obtain the $\lambda_{i,j,g}$, this is anyway unavoidable for downlink transmission and hence not part of the additional complexity resulting from the usage of the proposed technique. Further, note that since the Hessian of the objective function as shown in (18a) allows a block-diagonal structure, each iteration of the optimization problem involves solving a sparse linear system of equations which significantly reduces the overall complexity. If PRBG allocation is not considered, i.e., all the available PRBGs are allocated to every UE, the complexity is significantly reduced from ${\cal O}\left(\left(n_{RBG}n_{U}\right)^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$ to ${\cal O}\left(n_{U}^{3.5}\mathop{\mathrm{ln}}\left(1/\epsilon\right)\right)$, but this might come at the cost of throughput performance, especially in highly frequency selective channels. Further, we can always neglect the requirement on a certain error tolerance and fix the number of iterations to a predetermined number. This would further tradeoff performance for lower complexity.
@@
The computational burden of our method, requiring the execution of IPOPT every slot, presents a substantial challenge for real-time 6G implementation. To address this, we propose investigating both supervised and reinforcement learning approaches. Supervised learning could involve training a model on a dataset for which the labels have been obtained by the proposed technique, allowing for rapid prediction in subsequent slots. Alternatively, reinforcement learning, guided by our proposed technique as an expert policy, could uncover even better performing methods. Future research will focus on exploring these ML-based techniques, aiming to balance computational efficiency and performance.
@@ V Simulation Results
Figure 2: UL: Plots (with $90\%$ confidence interval) of the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot for the isolated cell case.
Figure 3: UL: Percentile ranks of the UE transmit power (left) and the PRB allocation for the isolated cell case.
Figure 4: UL: Plots (with $90\%$ confidence interval) of the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot with ICI.
Figure 5: UL: Percentile ranks of the UE transmit power (left) and the PRB allocation (right) for the case with ICI.
Figure 6: UL: Percentile ranks of the number of data layers per UE and the number of spatial layers per BS for the case with ICI.
Figure 7: DL: Plots (with $90\%$ confidence interval) of the UE GM rates (plain bar) and AM rates (hatched bars) in bits/slot.
Figure 8: DL: Percentile ranks of the number of data layers per UE and the number of spatial layers per BS.
Figure 9: DL: Percentile ranks of the BS transmit power (left) and the UE PRB allocation (right).
  1. 1.
  2. 2.
    Cells with ICI from the adjacent cells whose statistics vary from slot to slot depending on the paired UEs in the adjacent cells. In this case, there is no way of obtaining an accurate enough estimate of $\mathbf{R}_{g}$ (introduced in Section III-A) at the BS in order to accurately estimate the correct MCS for all the co-scheduled UEs unless the ICI is lower than or comparable to the thermal noise level.
For the uplink, we use QPSK, 16-QAM, 64-QAM, and 256-QAM while for the downlink, we also use 1024-QAM. Explicit data transmission, receiver equalization, and channel coding/decoding using 5G NR LDPC codes are performed. We assume the availability of perfect CSI but no open loop link-adaptation (OLLA) [17] or hybrid automatic repeat request (HARQ) whose inclusion, in practice, can be expected to make up for ill-effects of imperfect CSI. Therefore, MCS selection is done as detailed in [18] while taking $\mathbf{R}_{g}$ to be $\sigma^{2}\mathbf{I}$ in (3), where $\sigma^{2}$ is the thermal noise variance per PRB.
@@
$N_{UE}$ $8$ (maximum)
$n_{U}^{(i)}$ $4$ ($2$H×$\times 1$V×$\times 2$P), $\forall i$
@@
Number of cells
in the network $21$
Total number of UEs
in the network $210$
@@
Channel model 38.901 Urban Macro (UMa) NLoS
@@
UE OLPC parameters $P_{0}$ (dBm) $\in\{-85,-90,-100,-110\}$,
$(P_{0},\alpha)$ $\alpha\in\{0.85,1\}$
Channel coding 5G NR LDPC [15, Section 5.3.2]
with rate-matching
@@
MCS Table 5.1.3.1-2 (uplink),
Table 5.1.3.1-4 (downlink) of [8]
@@
HARQ No
Number of independent
UE drops 10
@@
Downlink precoder Block diagonalization ZF [1]
CSI aquistion frequency Once every $20$ slots
Number of iterations
of IPOPT $10$
V-A Uplink Results
Baseline power control: The total transmit power $P_{i}$ (dBm) for UE $i$ is obtained by setting the OLPC parameters $(P_{0},\alpha)$ in the following [19, Section 7]:
$P_{i}=\min\{P_{U,max},P_{0}+10\mathop{\mathrm{log}_{10}}\left(n_{PRB,i}\right)+\alpha PL_{i}\}$ (22)
@@ V-A Uplink Results
Baseline PRBG allocation: We assume that all the PRBGs are allocated to the UE unless power-limited. In the latter case, we assign a number of PRBGs (subject to a minimum of $4$) over which the UE has sufficient power to transmit, choosing the available PRBGs with the strongest channel gains for that UE.
@@
Fig. 2 shows the comparison (“BL" indicates “baseline") of the average GM and arithmetic mean (AM) UE rates, both computed per slot and then averaged across slots and independent UE drops. While the proposed technique is not targeted at maximizing the AM rates, they have been shown purely for illustration. Since there is no ICI, it is obvious that aggressively using as much of the available transmit power as required is generally the best strategy. Therefore, we use $r_{min}=0.23$ (corresponding to 4-QAM with the least coding rate in 5G NR) and $r_{max}=8$ (corresponding to the upper bound on the rate per data symbol for 256-QAM) for the proposed technique (Section III-A). While the proposed technique not only achieves the best GM UE rates, it also does so while requiring lower UE transmit powers than the naive, full transmit power scheme, as shown in Fig. 3.
For the case with ICI, since the transmit powers of the served UEs affect the reception in the neighbouring cells and the statistics of this ICI varies slot-to-slot, the best performing techniques usually choose the transmit powers conservatively so as to minimize the ICI. Otherwise, MCS selection would be significantly affected even for the baseline schemes. So, we conservatively take $r_{max}$ to be between $1$ and $2$. We also consider $\gamma=0.1$ for one of the baseline schemes so that a higher number of layers is selected for each UE. Fig. 4 highlights that the proposed scheme provides a significantly higher ($>50\%$) GM rates compared to the best baseline schemes while requiring a similar level of transmit powers per UE (for most baseline schemes) as shown in Fig. 5. The empirical cumulative distribution functions (CDFs) of the number of spatial layers per UE and the total number of spatial layers per BS are shown in Fig. 6. The trend clearly shows that while the AM rates can be improved by increasing $r_{max}$ at the cost of GM rates, the best balance is achieved for $r_{max}$ between $1.25$ and $1.75$.
@@ V-B Downlink Results
Baseline power allocation: Let $\mathbf{w}_{g,row}^{(k)}$ denote the $k^{th}$ row of $\mathbf{W}_{g}$, $k=1,\cdots,n_{B}$, $g=1,\cdots,n_{RBG}$, where $\mathbf{W}_{g}$ is the equivalent of $\mathbf{W}_{f,t}$ mentioned in Section II-B at a PRBG level. The columns of $\mathbf{W}_{g}$ are normalized to unity. Then, the precoder is taken to be $\mathbf{W}_{g}\leftarrow\sqrt{P}\mathbf{W}_{g}$ with $P=n_{RBG}P_{ant}/max_{k=1,\cdots,n_{B}}\left\{\sum_{g\in{\cal G}}n_{F}\|w_{g,row}^{(k)}\|^{2}\right\}$, where $n_{F}/n_{RBG}=12$, the number of subcarriers per PRBG.
Baseline PRBG allocation: For lack of a better alternative baseline scheme, we assume that all the PRBGs are allocated to all the UEs. Note that the techniques highlighted in [6][20] require reinforcement-learning and multiple PRB-looping which are out of scope of this paper.